Let A denote the area of the graph of f(x)=rootx on [0, 1] and let B denote the area of the graph of f(x)=x^2 on [0, 1]. Explain geometrically why A+B=1
1
Expert's answer
2020-05-14T17:49:08-0400

Let's consider the graph of "y = \sqrt{x}" for "x \in [0,1]."(green on the pic.2 and pic.1) Since "y = \sqrt{x}" , and "y \ge 0, x \ge 0" , we obtain that:

"y^2 = x".

Let's also notice, that there are two points of intersection of these functions, namely: "(0,0) , (1,1)" . Thus, graphs of given functions "\in" square with the side = 1. Its vertices are:"(0,0), (1,0), (1,1), (0,1)"


Thus if we consider "x" as a function and "y" as a variable, the area under the graph "x = y^2"(where y is a variable), which is equal to "S", according to the pic.2, is also equal(since these graphs are for same functions on same intervals) to the area of graph "y = x^2" (but with x as a variable)(NOT shaded part of not numbered picture).

So, according to the pic.2, A+S is the area of the square, mentioned above. But S is, correspondingly to the task, equal to B, and the area is obviously equal to "1 \cdot 1 = 1" .

Hence, A+B = 1.